A father is x years old and the son is y the sum of their age is 58. five years ago the father was five times as old as the son find the age of the two men
X+y=58
5(x-5y)=?
please help

Answer by stanbon(75887)

A father is x years old and the son is y.
the sum of their ages is 58.
---
x + y = 58
----
five years ago the father was five times as old as the son
(x-5) = 5(y-5)
----------------------
find the age of the two men
================================
x = 58-y
Substitute for "x" and solve for "y":
(58-y-5) = 5(y-5)
-y+53 = 5y-25
6y = 78
y = 13 yrs (son's age now)
---
Solve for "x"
x + y = 58
x + 13 = 58
x = 45 (father's age now)
---------------------------------------
Cheers,
Stan H.
==============